Suppose that ABC Ltd is considering purchasing one of three new processing machines. Either machine would make it possible for the company to produce its products more efficiently. Estimates regarding each machine are provided below:

Machine A Machine B Machine C
Original cost $79,000 $110,000 $244,000
Estimated life 7 years 8 years 10 years
Salvage value Nil Nil $30,000
Estimated annual cash inflows $30,000 $ 60,000 $58,500
Estimated annual cash outflows $ 7,000 $ 35,000 $18,500



If the projects cannot be repeated, which machine should ABC Ltd choose based on the NPV criteria at an 8% cost of capital? (9 marks)
1
Expert's answer
2020-10-28T12:38:49-0400

NPV = - I.O / (1+r)n + NCF /(1+r)n


a) "NPVa=-79,000+(30,000-7,000)\times\frac{(\frac{1}{1.08})^7-1}{\frac{1}{1.08}-1}=50,326.23"


"NPVb=-110,000+(60,000-35,000)\times\frac{(\frac{1}{1.08})^8-1}{\frac{1}{1.08}-1}=45,159.25"


"NPVc=-244,000+(58,500-18,500)\times\frac{(\frac{1}{1.08})^{10}-1}{\frac{1}{1.08}-1}-\frac{30,000}{1.08^{10}}=31,979.72"


 ABC Ltd should choose project A since it has the highest NPV